OpenStudy (anonymous):

draw a spinner with two sections red, blue such that the probability of the event of getting (blue,blue) on two spins is 25/36

OpenStudy (anonymous):

the probability of getting blue blue would be \(\frac{5}{6}\) because \(\frac{5}{6}\times \frac{5}{6}=\frac{25}{36}\)

OpenStudy (anonymous):

so make sure that the blue section is \(\frac{5}{6}\) the total area, red section only \(\frac{1}{6}\)

OpenStudy (anonymous):

 something like this

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw